OpenStudy (anonymous):

Simplify. Please show all of your work 5-2^4 /16.3 -4

OpenStudy (anonymous):

\[(5-2^{2}) \div (16.3\times-4)\] you meant ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a) \[(5−2^{2})÷(16.3×−4)\] 1 ÷ 4.075 b) if you meant by \[[(5−2)^{2}÷(16.3×−4)\] than 9 ÷ 4.075

OpenStudy (anonymous):

what if the 16.3-4 was under the 5-2^4 would it be deterrent

OpenStudy (anonymous):

different

OpenStudy (anonymous):

\[\left(\begin{matrix}5-2^{2} \\ 16.3-4\end{matrix}\right)\] is the same thing as (5−2)2÷(16.3×−4) unless you mean the negative four was a subtract four from the 16.3

OpenStudy (anonymous):

is it 16.3 -4 or (16.3x-4) ?

OpenStudy (anonymous):

just 16.3-4

OpenStudy (anonymous):

ohh okay, then \[\left(\begin{matrix}5-2^{^{2}} \\ 16.3 - 4\end{matrix}\right)\] or \[(5-2)^{2} \div (16.3 - 4 ) \] is the most likely answer.

OpenStudy (anonymous):

the one before was good then

OpenStudy (anonymous):

yesh xD